OpenStudy (anonymous):

Please help me with solving an equation! I'll fan and medal 8-2b=-2/3(12 b+6)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I need an explanation

OpenStudy (anonymous):

and it's not 3, it's -2

OpenStudy (anonymous):

(I just don't know how to get that answer)

OpenStudy (butterflydreamer):

\[8 - 2b = -\frac{ 2 }{ 3 } (12b+6)\] The first thing you want to do is to get rid of the fraction :) So multiply both sides by 3. Like this: \[3 ( 8 - 2b) = 3 \times -\frac{ 2 }{ 3 } (12b+6)\]

OpenStudy (butterflydreamer):

now what would you get when... \[3 \times -\frac{ 2 }{ 3 } = ?\]

OpenStudy (anonymous):

-2?

OpenStudy (butterflydreamer):

excellent! So now we know this: \[3 (8-2b) = -2 (12b + 6)\] Now we want to get rid of the brackets so expand the brackets

OpenStudy (anonymous):

(12b + -2) x (6+-2)? i'm not sure

OpenStudy (butterflydreamer):

hmm...not quite. OKay, let's just focus on everything on the LEFT side of the '=' sign. \[3 (8 - 2b)\] So to expand this, we multiply everything in the brackets by 3. \[3 ( 8 - 2b) = (3 \times 8 ) - (3 \times 2b)= ?\]

OpenStudy (butterflydreamer):

what does 3 times 8 = ? And what does 3 times 2b = ?

OpenStudy (anonymous):

24 and 6b?

OpenStudy (butterflydreamer):

yeeess! So then, \[3 ( 8 - 2b) = 24 - 6b\] So that's the left side covered of the equation: \[3( 8 - 2b) = -2 ( 12b + 6)\] Now if we look at the right side of the = sign, we have: \[-2 ( 12b + 6) = ?\] Can you expand this one (think of it like -2 times 12b PLUS -2 times 6) ?

OpenStudy (anonymous):

(-2 x 12b) + (-2 x 6)

OpenStudy (butterflydreamer):

yepp and what would that equal to? -2 x 12b = ? -2 x 6 = ?

OpenStudy (anonymous):

-24b and -12? I have a question too, does it matter if the terms are unlike when adding and subtracting/ multiplying and dividing?

OpenStudy (butterflydreamer):

that's correct! So now we know, \[-2 ( 12b + 6) = -24b - 12\] So if we put it all together, this is what we have done: \[8 - 2b = \frac{ -2 }{ 3 } (12b + 6)\] We wanted to "get rid of the fraction" so we multiplied both sides by 3 \[3 ( 8 - 2b) = 3 \times - \frac{ 2 }{ 3 } (12b + 6)\] \[3 ( 8 - 2b) = -2 (12b + 6)\] Then we expanded the brackets or "multiplied everything out" \[24 - 6b = -24b - 12\] Now from here, you just want to collect like terms :)

OpenStudy (anonymous):

Ok so 6b-24b-24-12?

OpenStudy (butterflydreamer):

what happened to your equal sign? \[24 - 6b = -24b - 12\] Add 24b to both sides: \[24 - 6b + 24b = -24b - 12 + 24b\] Simplify & collect like terms

OpenStudy (butterflydreamer):

so, 24 - 6b + 24b = ? and then -24b - 12 + 24b = ?

OpenStudy (anonymous):

Are we adding -24b on both sides?

OpenStudy (butterflydreamer):

we are adding +24b on both sides (so this way, we can get rid of the -24b on the right side)

OpenStudy (anonymous):

so 36b?

OpenStudy (butterflydreamer):

what does 24 - 6b + 24b = ?

OpenStudy (butterflydreamer):

e.g. 24 - 6bananas + 24bananas = ? (you want to collect like terms so only the bananas can be grouped together). 24 - 6bananas + 24bananas = 24 + 18bananas right? so the same thing applies here. 24 - 6b + 24b = 24 + 18b

OpenStudy (butterflydreamer):

now, what does -24b - 12 + 24b = ? You can rearrange this equation if it is easier for you: 24b - 24b - 12 = ?

OpenStudy (anonymous):

-12?

OpenStudy (butterflydreamer):

YES! :DSo now, 24 - 6b = -24b - 12 Then we added +24b to both sides: 24 - 6b + 24b = -24b - 12 + 24b We collected like terms: 24 + 18b = -12 The next step is to isolate the 18b because we want to find out what "b" equals to right? (We want something that looks like 18b =...) SO we need to SUBTRACT 24 from both sides. 24 + 18b - 24 = -12 - 24 18b = -12 - 24 What does -12 - 24 = ?

OpenStudy (anonymous):

-36?

OpenStudy (butterflydreamer):

yaaaa. So, 18b = -12 - 24 18b = - 36 We have 18b but we only need to find out what "b" equals to. Divide both sides by 18 and you're done!

OpenStudy (butterflydreamer):